# Linear Representations of Political Perspective Emerge in Large Language Models

### v1

@@ 2 Preliminaries
where $\bm{u}_{v}\in\mathbb{R}^{D}$ is the unembedding of possible token value $v$, and $\bm{r}_{L}\in\mathbb{R}^{D}$ is the final vector in the transformer’s “residual stream” (Elhage et al., 2021), which evolves over layers $\ell=1,\dots,L$ as:
@@ Results
We also provide results for the ensembled models in Table A3), where we find that performance tapers around $K=32$, at which $\widehat{\rho}^{\textrm{CV}}_{K}$ is 0.87 for Llama-2-7b-chat, 0.864 for Mistral-7b-instruct, and 0.885 for Vicuna-7b. In Figure 3 we also plot the ensembled predictions for Mistral-7b-instruct and highlight examples of well-known lawmakers; the same plot for all three models is in Figure 1(c) and Figure 2(c).
@@
We do not observe substantial improvements when using such non-linear probes. For Llama-2-7b-chat, the most predictive linear probe had a cross-validation Spearman correlation of 0.858 while the best non-linear probe achieved 0.856. For Vicuna-7b, the difference was larger, with the linear and non-linear probes achieving 0.861 and 0.872, respectively. But for Mistral-7b-instruct, the order was reversed, with the linear and non-linear probes achieving 0.846 and 0.841, respectively. These results support the linear representation hypothesis for political ideology in the sense that linear functions of certain attention heads predict DW-NOMINATE approximately as well as non-linear functions of any others.
@@
The results are given in Figure 3(d) and Table A4. The probes trained on randomly permuted labels provide a baseline Spearman correlation of around 0.15, representing chance performance. Probes trained to predict the non-monotonic transformation perform poorly, with the best-performing heads achieving correlations of around 0.5. As might be expected, probes trained to predict the cubic transformation do much better, with the best-performing heads achieving rank correlations close to 0.83. In addition to rank correlation, which should not be sensitive to monotonic transformations, we also include in Table A4 the cross-validation $R^{2}$ values of the different probes. These tell a different story, with the cubic probes exhibiting values of around 0.6 compared to values of 0.8 achieved by the original.
@@
Figure 3: Ideological perspectives of U.S. lawmakers and news media as predicted by the activations of the $K=32$ most predictive attention heads of Llama-2-7b-chat. Negative values correspond to left-leaning perspectives, while positive values correspond to right-leaning perspectives. The $x$-axis represents the predicted political slant ($\widehat{y}_{K=32}^{{(i)}}$) for each entity (i.e., lawmakers or news media). The $y$-axis represents the formerly validated ideological scores (DW-NOMINATE or Ad Fontes Media scores). See Figures 1(c) and 2(c) for the complete results across all models.
@@ 6 Trained Probes Can Be Used to Steer Political Perspective
To measure the political perspective of each generated essay, we first recruited 10 human annotators from CloudResearch Survey—three Democrats, four independents, and three Republicans—and had them rate a random sample of the essays on a 7-point scale from “Strongly conservative” to “Strongly liberal”. We then instructed GPT-4o (gpt-4o-2024-08-06) to rate the same essays on the same scale (see Section A.3 for the exact prompt) and measured the inter-rater reliability between the GPT ratings and the average human ratings. GPT-4o’s ratings were very close to the humans’, with an intra-class correlation of 0.91, which we considered license to use it for rating the remaining essays; see Section A.4 for more details.
@@ Results
Llama-2-7b-chat displayed the highest correlation of 0.609 between $\alpha$ and political slant, followed by Mistral-7b-instruct at 0.405, and Vicuna-7b at 0.394. Political slant increased steadily as $\alpha$ increased, particularly in Llama-2-7b-chat, suggesting that this model is more sensitive to intervention. We also experimented with intervening on different numbers of attention heads $K$, and found that intervening on more led to greater effectiveness; see Figure A5.
@@
One might wonder whether the interventions we describe will continue to be effective at steering when discussing policy issues not described in the model’s pre-training data. In Section A.6, we show that interventions remained effective when models were instructed to write about two events that fell after Llama-2-7b-chat’s cutoff: 1) the U.S. ADVANCE Act, and 2) the 2023 United Auto Workers (UAW) Strike.
@@ 8 Conclusions, Limitations, and Future Directions
Our study has several limitations. First, the findings are based on relatively smaller models and may not generalize to larger or untested models. Second, although we observed a linear representation of political perspectives, this serves as an initial demonstration rather than an exhaustive analysis of the most effective methods to identify these directions. Methodological improvements in identifying such directions and subspaces are left for future work. Third, our research is U.S. centric and may not generalize to less polarized political environments, where linear representations of ideologies may be less effective representations (See Section A.10 for details). In such settings, however, we may characterize ideologies as the simplex of more than two “archetypal” or extreme political perspectives (Seth & Eugster, 2016). Fourth, we use GPT-4o to evaluate political slant; however, there is potential for bias when using an LLM as an evaluator. Although we validate GPT-4o’s evaluations against politically balanced human annotators, we recommend that future research using our methods continue to validate LLM-generated annotations against human annotations to triangulate and mitigate any inherent biases. Finally, future research could explore whether political ideology can be combined with other features in a meaningful, linear fashion or show how this linearity generalizes to other tasks or representations beyond political ideology. Other dimensions of cultural perspective (e.g., social class, gender) (Kozlowski et al., 2019) or knowledge and experience-based expertise were not explored in this paper. We hope that future research will investigate this promising direction and its potential to craft and steer customizable LLM agent perspectives.
@@
The data and code for reproducing our results are available on Github [^5] https://github.com/JunsolKim/RepresentationPoliticalLLM. All experiments were conducted on an internal computing cluster equipped with: AMD EPYC 7313 16-Core Processor, NVIDIA A100 80GB, 100GB of memory, and 100GB of disk space.
@@ A.1 Model Overview
  • Llama-2-7b-chat: This model is part of the Llama-2 family, developed by Meta, with 7 billion parameters. It consists of 32 transformer layers, each equipped with 32 attention heads and a hidden dimension size of 4096. The model is optimized for conversational tasks through fine-tuning.
  • Mistral-7b-instruct-v0.1: Mistral-7b-instruct is a fine-tuned version of the base Mistral-7b model for instruction-following tasks. Similar to Llama-2-7b-chat, Mistral-7b-instruct-v0.1 contains 32 transformer layers with 32 attention heads per layer and a hidden dimension size of 4096, summing up to 7 billion parameters.
  • Vicuna-7b-v1.5: Vicuna-7b is a fine-tuned version of Llama-2, optimized for chatbot tasks. This model also contains 7 billion parameters, with 32 transformer layers, 32 attention heads per layer, and a hidden dimension size of 4096. The fine-tuning focuses on generating dialogue responses.
@@ A.2 Trained Probes Detect Political Perspective Token-by-Token
As shown in Figure A6, even though we do not explicitly prompt LLMs to adopt a political stance, we find that the models spontaneously generate texts reflecting either left- or right-leaning perspectives, as indicated by the token-level values of political slant $\widehat{y}^{{(i,t)}}_{\ell,h}$. Specifically, Mistral-7b-instruct shows an average $\widehat{y}^{{(i,t)}}_{\ell,h}$ of 0.143, exhibiting a slightly more right-leaning slant compared to Llama-2-7b-chat (0.054) and Vicuna-7b (0.048). Interestingly, Mistral-7b-instruct also displays about twice the diversity in perspectives, as measured by the standard deviation of $\widehat{y}^{{(i,t)}}_{\ell,h}$, with $\sigma=0.648$ compared to Llama-2-7b-chat ($\sigma=0.349$) and Vicuna-7b ($\sigma=0.356$).
@@ A.10 Cross-National Analysis of Political Slant Representations
We utilize the following prompt to simulate thideological perspectives of lawmakers in these parties:
@@
A.11 Replication on Gemma-2-9b
Some might question whether political ideologies are similarly represented in the middle layers of LLMs outside the Llama family. We successfully replicated our analysis on the Gemma-2-9b model and found that it also exhibits a linear representation of ideological slant in its middle layers. See Figure A10 for details.
Table A1: Effect of regularization parameter $\lambda$ on probe performance. We compare $\widehat{\rho}^{\textrm{CV}}_{\ell,h}$ at the most predictive attention head across models for different $\lambda$ values. The value $\lambda=1$ optimizes these metrics across the three models.
@@ A.11 Replication on Gemma-2-9b
Llama2-7b-chat 0.854 0.152 0.550 0.842
Mistral-7b-instruct 0.846 0.140 0.533 0.841
@@
Figure A10: Predictive performance of linear probes for each attention head across all layers in gemma-2-9b. Performance is measured using Spearman rank correlation, with darker shades indicating stronger correlations.

### v2

@@ 2 Preliminaries
where $\bm{u}_{v}\in\mathbb{R}^{D}$ is the unembedding of possible token $v$, and $\bm{r}_{L}\in\mathbb{R}^{D}$ is the final vector in the transformer’s “residual stream” (Elhage et al., 2021), which evolves over layers $\ell=1,\dots,L$ as:
@@ Results
We also provide results for the ensembled models in Table A3, where we find that performance tapers around $K=32$, at which $\widehat{\rho}^{\textrm{CV}}_{K}$ is 0.87 for Llama-2-7b-chat, 0.864 for Mistral-7b-instruct, and 0.885 for Vicuna-7b. In Figure 3 we also plot the ensembled predictions for Llama-2-7b-chat and highlight examples of well-known lawmakers; the same plot for all three models is in Figure 1(c) and Figure 2(c).
@@
We do not observe substantial improvements when using such non-linear probes. For Llama-2-7b-chat, the most predictive linear probe had a cross-validation Spearman correlation of 0.854 while the best non-linear probe achieved 0.855. For Vicuna-7b, the difference was larger, with the linear and non-linear probes achieving 0.861 and 0.872, respectively. But for Mistral-7b-instruct, the order was reversed, with the linear and non-linear probes achieving 0.846 and 0.838, respectively. These results support the linear representation hypothesis for political ideology in the sense that linear functions of certain attention heads predict DW-NOMINATE approximately as well as non-linear functions of any others.
@@
The results are given in Figure 3(d) and Table A4. The probes trained on randomly permuted labels provide a baseline Spearman correlation of around 0.15, representing chance performance. Probes trained to predict the non-monotonic transformation perform poorly, with the best-performing heads achieving correlations of around 0.5. As might be expected, probes trained to predict the cubic transformation do much better, with the best-performing heads achieving rank correlations close to 0.84. In addition to rank correlation, which should not be sensitive to monotonic transformations, we also include in Table A4 the cross-validation $R^{2}$ values of the different probes. These tell a different story, with the cubic probes exhibiting values of around 0.6 compared to values of 0.8 achieved by the original.
@@
Figure 3: Ideological perspectives of U.S. lawmakers and news media as predicted by the activations of the $K=32$ most predictive attention heads of Llama-2-7b-chat. Negative values correspond to left-leaning perspectives, while positive values correspond to right-leaning perspectives. The $x$-axis represents the predicted political slant ($\widehat{y}_{K=32}^{{(i)}}$) for each entity (i.e., lawmakers or news media). The $y$-axis represents the previously validated ideological scores (DW-NOMINATE or Ad Fontes Media scores). See Figures 1(c) and 2(c) for the complete results across all models.
@@ 6 Trained Probes Can Be Used to Steer Political Perspective
To measure the political perspective of each generated essay, we first recruited 10 human annotators from CloudResearch Survey—three Democrats, four independents, and three Republicans—and had them rate a random sample of the essays on a 7-point scale from “Strongly conservative” to “Strongly liberal”. We then instructed GPT-4o (gpt-4o-2024-08-06) to rate the same essays on the same scale (see Section A.3 for the exact prompt) and measured the inter-rater reliability between the GPT ratings and the average human ratings. GPT-4o’s ratings were very close to the humans’, with an intraclass correlation of 0.91, which we considered license to use it for rating the entire essays; see Section A.4 for more details.
@@ Results
When $K\in\{64,80,96\}$, Llama-2-7b-chat displayed the highest correlation of 0.607 between $\alpha$ and political slant, followed by Mistral-7b-instruct at 0.396, and Vicuna-7b at 0.381. Political slant increased steadily as $\alpha$ increased, particularly in Llama-2-7b-chat, suggesting that this model is more sensitive to intervention. We also experimented with intervening on different numbers of attention heads $K$, and found that intervening on more led to greater effectiveness; see Figure A5.
@@
One might wonder whether the interventions we describe will continue to be effective at steering when discussing policy issues not described in the model’s pre-training data. In Section A.6, we show that interventions remained effective when models were instructed to write about two events that fell after Llama-2-7b-chat’s pre-training cutoff: 1) the U.S. ADVANCE Act, and 2) the 2023 United Auto Workers (UAW) Strike.
@@ 8 Conclusions, Limitations, and Future Directions
Our study has several limitations. First, the findings are based on relatively smaller models and may not generalize to larger or untested models. Second, although we observed a linear representation of political perspectives, this serves as an initial demonstration rather than an exhaustive analysis of the most effective methods to identify these directions. Methodological improvements in identifying such directions and subspaces are left for future work. Third, our research is U.S. centric and may not generalize to less polarized political environments, where linear representations of ideologies may be less effective representations (See Section A.10 for details). In such settings, however, we may characterize ideologies as the simplex of more than two “archetypal” or extreme political perspectives (Seth & Eugster, 2016). Fourth, we use GPT-4o to evaluate political slant; however, there is potential for bias when using an LLM as an evaluator. Although we validate GPT-4o’s evaluations against politically balanced human annotators, we recommend that future research using our methods continue to validate LLM-generated annotations against human annotations to triangulate and mitigate any inherent biases. Future research could also explore whether there are linear representations of more granular or intersected forms of political ideology. Other dimensions of cultural perspective (e.g., social class, gender) (Kozlowski et al., 2019) or knowledge and experience-based expertise were not explored in this paper. We hope that future research will investigate this promising direction and its potential to craft and steer customizable LLM agent perspectives.
@@
The data and code for reproducing our results are available on Github [^5] https://github.com/JunsolKim/RepresentationPoliticalLLM.
@@ A.1 Model Overview
  • Llama-2-7b-chat: This model is part of the Llama-2 family, developed by Meta, with 7 billion parameters. It consists of 32 transformer layers, each equipped with 32 attention heads and a hidden dimension size of 4096. The model is optimized for conversational tasks.
  • Mistral-7b-instruct-v0.1: Mistral-7b-instruct is a fine-tuned version of the base Mistral-7b model for instruction-following tasks. Similar to Llama-2-7b-chat, Mistral-7b-instruct-v0.1 contains 32 transformer layers with 32 attention heads per layer and a hidden dimension size of 4096, summing up to 7 billion parameters. The model is optimized for conversational tasks.
  • Vicuna-7b-v1.5: Vicuna-7b is a fine-tuned version of Llama-2, optimized for conversation tasks. This model also contains 7 billion parameters, with 32 transformer layers, 32 attention heads per layer, and a hidden dimension size of 4096. The fine-tuning focuses on generating dialogue responses.
@@ A.2 Trained Probes Detect Political Perspective Token-by-Token
As shown in Figure A6, even though we do not explicitly prompt LLMs to adopt a political stance, we find that the models spontaneously generate texts reflecting either left- or right-leaning perspectives, as indicated by the token-level values of political slant $\widehat{y}^{{(i,t)}}_{\ell,h}$. Specifically, Mistral-7b-instruct shows an average $\widehat{y}^{{(i,t)}}_{\ell,h}$ of 0.143, exhibiting a slightly more right-leaning slant compared to Llama-2-7b-chat (0.054) and Vicuna-7b (0.048). Interestingly, Mistral-7b-instruct also displays about twice the diversity in perspectives, as measured by the standard deviation of $\widehat{y}^{{(i,t)}}_{\ell,h}$, with $\sigma=0.649$ compared to Llama-2-7b-chat ($\sigma=0.349$) and Vicuna-7b ($\sigma=0.356$).
@@ A.10 Cross-National Analysis of Political Slant Representations
We utilize the following prompt to simulate the ideological perspectives of lawmakers in these parties:
@@
A.11 Replication on Gemma-2-2b
Some might question whether political ideologies are similarly represented in the middle layers of LLMs outside the Llama family. We successfully replicated our analysis on the Gemma-2-2b model and found that it also exhibits a linear representation of ideological slant in its middle layers. See Figure A10 for details.
Table A1: Effect of regularization parameter $\lambda$ on probe performance. We compare $\widehat{\rho}^{\textrm{CV}}_{\ell,h}$ at the most predictive attention head across models for different $\lambda$ values. The value $\lambda=1$ optimizes these metrics across the three models, except for Vicuna-7b.
@@ A.11 Replication on Gemma-2-2b
Llama2-7b-chat 0.854 0.152 0.551 0.842
Mistral-7b-instruct 0.846 0.140 0.534 0.841
@@
Figure A10: Predictive performance of linear probes for each attention head across all layers in gemma-2-2b. Performance is measured using Spearman rank correlation, with darker shades indicating stronger correlations.